Carland, Inc., has a project available with the following cash flows. If the required return for the project is 9.6 percent, what is the project's NPV? Year Cash Flow 0 −$275,000 1 86,700 2 109,100 3 126,300 4 77,700 5 −13,700
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=86,700/1.096+109,100/1.096^2+126,300/1.096^3+77,700/1.096^4

=$319713.42

Present value of outflows=275,000+13,700/1.096^5

=283662.99

NPV=Present value of inflows-Present value of outflows

=319713.42-283662.99

=$36050.43(Approx).